Gay-Lussacs Law: Ti T Dont forget to convert C to K obni 1. 9.0L of a gas is exerts a pressure of 83.0 kPa at 35.0°C. What would be the required temperature (in ℃elsius) to change the pressure to standard pressure? 4
Answer #1

P1; 83 kPa

Standard Pressure (P2); 101.3kPa

T1= 35 C

= 308.15 K (Kelvin)

P1/T1 = P2/T2

83 kPa/308.15 K =   308.15 K/T2

0.269 = 308.15 K/T2

T2 = 308.15/0.269

T2= 1145.53903 K